OpenStudy (anonymous):

Y varies inversely as twice x when x=9, y=2, find y when x=3? can someone work this out for me and then explain how they did it?

OpenStudy (lgbasallote):

do you know the equation for inverse variation?

OpenStudy (anonymous):

no this was all i was given...:/

OpenStudy (lgbasallote):

it's never given...it's memorized... anyway the equation for the inverse variation is \[y = \frac kx\] okay?

OpenStudy (anonymous):

okay yes got that...

OpenStudy (lgbasallote):

good. here, it says y varies inversely as TWICE x so your equation would be \[y = \frac k{2x}\] follow so far?

OpenStudy (anonymous):

yes i follow

OpenStudy (lgbasallote):

now it says x=9 when y = 2 so substitute x = 9 and y = 2 into \[y = \frac k{2x}\] then solve for k

OpenStudy (anonymous):

so now it would be 2= k/2(2) Right now i solve for k

OpenStudy (lgbasallote):

no 2 = k/2(9)

OpenStudy (anonymous):

oppies my bad

OpenStudy (lgbasallote):

now solve for k

OpenStudy (anonymous):

okay so i have 2=k/2(9) now i multiply 2(9)=18 then divide by 2 and k =9 2=k/18 k=9

OpenStudy (anonymous):

i thinkk???

OpenStudy (anonymous):

right ?? @igbasallote

OpenStudy (lgbasallote):

you don't divide by 2..you multiply both sides by 18